# Learning Fair Representations with Kolmogorov-Arnold Networks

### v1

@@ Introduction
On training the KAN models on $D_{(1)}$ dataset, we observe the Adam-based model to offer stronger predictive power ($0.45\%$ incr. in Acc., $3.41\%$ incr. in AUROC) compared to OAdam, which enhances fairness outcomes significantly. However, the model trained using ADOPT exhibits a balanced performance across fairness and accuracy. On the other hand, training the KAN models on the larger $D_{(2)}$ dataset yields notable improvements in predictive performance and fairness. Although all setups trained on $D_{(2)}$ show significant performance improvement, the ADOPT-based model stands out by achieving balanced optimization between accuracy ($0.15\%$, and $0.82\%$ incr. in Acc. compared to Adam setup) and fairness metrics ($\{15.14\%,18.48\%\}$ incr. in $p\%$-Rule compared to Adam setup). Meanwhile, the $B_{(1)}$ and $B_{(1)}^{\prime}$ models under-perform relative to the KAN models, especially on the $D_{(2)}$ dataset. This suggests KAN-based adversarial learning to be more effective in maintaining fairness-accuracy trade-offs while capturing complex patterns, even in a feature-rich setting. State-of-the-art baseline models, $B_{(2)}$ and $B_{(3)}$, display varying behavior. $B_{(2)}$ achieves high fairness on $D_{(1)}$ but suffers a drop on $D_{(2)}$, whereas $B_{(3)}$ starts with lower fairness on $D_{(1)}$ but improves substantially on $D_{(2)}$ with proper fine-tuning.
@@
The empirical evidence demonstrates that the KAN-based adversarial debiasing framework, regardless of the choice of optimizer, consistently outperforms the baseline models across both the datasets. This can be largely attributed to the adaptive policy usage within the training module. Its ability to dynamically adjust and bound the fairness constraint, $\lambda$, achieves to secure a balanced performance. Amongst all the configurations, the model trained on ADOPT demonstrates comparative superior balanced performance. This observation suggests ADOPT to be well suited to the spline-based architecture of KANs, owing to their adaptive convergence behavior (taniguchi2024adopt).
@@
Figure 2: Ablation study to compare predictive performance and fairness scores across proposed frameworks with different spline knot complexities, and baseline models under varying optimization strategies, evaluated on dataset $D_{(1)}$. (a) Illustrates metrics of the proposed KAN-based adversarial frameworks using three different optimizer techniques (Adam- $o_{1}$, OAdam- $o_{2}$, ADOPT- $o_{3}$). Similarly, (b) depicts the Baseline model performance across varying optimizers.
@@
To investigate how KAN spline knot complexity, $k$, affects the adversarial learning process, we perform an ablation study focusing on three spline orders $k=\{3,4,5\}$. We chose the $D_{(1)}$ dataset for the experiments due to its controlled small-scale setting. It enables focused assessment of the model’s sensitivity to architectural changes under constrained data conditions. In addition to the KAN-based adversarial learning models, $K_{(1)}$, we also train the baseline model, $B_{(1)}$, across the three optimizer settings, and the state-of-the-art baseline models, $B_{(2)}$, and $B_{(3)}$ across the Adam optimizer. As illustrated in Figure 2(a), increasing the spline order $k$ leads to subtle variations in accuracy and fairness metrics. Although the $K_{(1)}$ models exhibit relatively stable accuracy across different $k$ setups, the fairness metrics, measured via the $p\%\text{-Rule}$, tend to improve with higher spline orders. This suggests that more expressive spline functions, in an adversarial setting, can better model fair representations. Although higher $k$ (e.g., $k=4,5$) may introduce convergence instability, as observed in Figure 3, the model trained with ADOPT exhibits comparatively stable training dynamics than those trained on Adam or OAdam. Next, when comparing the $K_{(1)}$ models with the baseline MLP models, $B_{(1)}$, we observe that the former consistently outperforms the latter in achieving a balance between accuracy and fairness. As observed in Figure 2(b), $B_{(1)}$ models have a higher fairness score but comparatively lower accuracy across the three optimizer settings. This further empirically substantiates the efficiency of the $\lambda$-penalty update policy utilized in our proposed $K_{(1)}$ models. The state-of-the-art baseline models, $B_{(2)}$ and $B_{(3)}$, show strong performance in fairness but incur slight drops in accuracy as seen in Figure 2(b). Although the significant spikes in the $K_{(1)}$ model training loss convergence plot, as seen in Fig. 3, can be attributed to the smoothness-dimension trade-off (samadi2024smooth) in KANs. This induces oscillatory behavior during training, but is inherent to preserve universal approximation. A more formal discussion of this property is provided in the Appendix.
Nonetheless, our findings help us empirically establish the efficiency of our proposed adversarial learning framework despite the fluctuations in KAN-based setups.
@@
In this study, we propose a KAN-based adversarial learning framework with adaptive $\lambda$ policy to mitigate socioeconomic bias in a college admission setting. We discuss the various structural properties of KANs, providing proper mathematical justification behind the adversarial robustness of the networks. Through extensive experiments on two distinct sets of real-world admissions datasets, we demonstrate the efficacy of our proposed framework by consistently outperforming standard baselines while achieving a balance between the inherent accuracy-fairness tradeoff. While our current proposed framework effectively mitigates output-level bias, it opens up several promising enhancements. The incorporation of explicit feature-level bias detection through a reweighting mechanism could improve the overall fairness of the model while also helping to adhere to any data privacy constraints. Additionally, extending the current framework by incorporating alternative fairness definitions beyond DP could provide new insights. We believe this work offers a promising step towards aligning algorithmic performance with social responsibility, encouraging future explorations into KAN-based fairness-aware systems that uphold both ethical and performance standards in critical decision-making.
@@ Appendix A Additional Theoretical Analysis
As provided in Eq. 6, $f$ is composed of inner functions $\phi$ (=h(.)=h(.)), and outer functions, $\psi$, where each of them is a smooth spline. Let the inner sum be defined as , $v_{k}(x)=\sum_{i=1}^{n}\phi_{k_{i}}(x_{i})$ Then we can write $f$ as $f(x_{1},\dots,x_{n})=\sum_{k=1}^{2n+1}\psi_{k}(v_{k}(x))$.
@@
$\left|\frac{\partial z_{k}}{\partial x_{j}}\right|=\left|\phi_{k_{j}}^{\prime}(x_{j})\right|\leq B$
@@
$\frac{\partial f}{\partial x_{j}}=\sum_{k=1}^{2n+1}\psi_{k}^{\prime}\left(z_{k}(x)\right)\cdot\phi_{k_{j}}^{\prime}(x_{j})$
@@
The input $x_{i}$ is summed at every neuron in a KAN structure, and then a univariate spline $\psi(x)$ is applied. Let the sum of the inner functions be denoted as $u$. Both $\phi$, and $\psi$ are twice continuously differentiable with bounded second order derivatives. Then, we can write, $|\phi_{k_{i}}^{\prime\prime}|\leq\beta_{1}$, $|\psi_{k}^{\prime\prime}|\leq\beta_{2}$, and the derivative of inner sum $u^{\prime}(x)$ is bounded by some constant $B>0$. Using chain rule, the second derivative of the composite function $f(x)=\psi_{k}(u(x))$ can be expressed as,
$f^{\prime\prime}(x)=\psi_{k}^{\prime\prime}(u(x))\cdot[u^{\prime}(x)]^{2}+\psi_{k}^{\prime}(u(x))\cdot u^{\prime\prime}(x)$
@@
$|f^{\prime\prime}(x)|\leq\beta_{2}B^{2}+|\psi_{k}^{\prime}(u(x))|\cdot\beta_{u}$
where, $\beta_{u}$ is an upper bound on $u^{\prime\prime}(x)$. Thus each composition $\psi_{k}(\sum_{i}\phi_{k_{i}}(x_{i}))$ is $\beta$-smooth. More generally, the entire KAN, $f$, is a nested composition of such units. Hence, if each layer’s activation function is twice continuously differentiable, then the final composed function also stays $C^{2}$ smooth (deboor1978splines).
@@
In this appendix section, we discuss a supplementary special case for analyzing the fairness bounds in an adversarial debiasing setup. Although this is not a part of our main algorithmic contribution, but serves to provide additional insight into the behavior of the framework under the following conditions.
We constraint the adversary model, $g$, through gradient penalty in order to satisfy and hold the following case (arjovsky2017wasserstein). Let $(X,d_{\mathcal{X}})$ and $(Y,d_{\mathcal{Y}})$ be the metric spaces. Given, the KAN classifier function $f:X\to Y$ is $L$-Lipschitz, i.e. $d_{\mathcal{Y}}(f(x),f(y))\leq L\cdot d_{\mathcal{X}}(x,y)$ for all $x,y\in X$.
@@
For any distribution $\mu$, and $\nu$ on $X$ function $\varphi$,
[…]
where $\varphi||_{\text{Lip}}\leq 1$ (villani2008optimal).
@@
Let $\mu_{0}$, $\mu_{1}$ be two distributions on $X$, and let $f:X\to Y$ be $L$-Lipschitz, then
@@
$=L\,d_{\mathcal{X}}(x,y)$ (7)
Let $\varphi=\frac{1}{L}g\circ f$. Then $\varphi$ is 1-Lipschitz on $X$ and by Definition B,
$\mathbb{E}_{\mu_{0}}[g\circ f]-\mathbb{E}_{\mu_{1}}[g\circ f]=L(\mathbb{E}_{\mu_{0}}[\varphi]-\mathbb{E}_{\mu_{1}}[\varphi])\leq LW_{1}(\mu_{0},\mu_{1})$
@@
[…]
@@
Let $Z\in\{0,1\}$ be a sensitive attribute, then, $\mu_{0}:=\mathbb{P}(X|Z=0)$, $\mu_{1}:=\mathbb{P}(X|Z=1)$. For any KAN, $f$, with Lipschitz constant $L_{f}$, in order to obtain a 1-Lipschitz adversary $g$, the empirical fairness objective can be stated as,
$L_{adv}(g,f):=\mathbb{E}[g(f(X))|Z=0]-\mathbb{E}[g(f(X))|Z=1]$
that satisfies $|L_{adv}(g,f)|\leq L_{f}\cdot W_{1}(\mu_{0},\mu_{1})$. Hence, maximizing $L_{adv}(g,f)$ over 1-Lipschitz $g$ yields $W_{1}(f_{\#}\mu_{0},f_{\#}\mu_{1})$.
@@ Appendix C Additional Experimental Results
All the experiments, excluding those involving $D_{(2)}$ datasets, have been conducted on MacBook Air with Apple M3 chip, and 24GB unified memory, running macOS. Large scale training on the $D_{(2)}$ datasets were performed on a remote Linux server with four NVIDIA RTX A6000 GPUs (49GB each), using CUDA 12.4. We have used Python-based NumPy and PyTorch packages to train the models.
@@
Table 2: Fairness metrics for the KAN-based adversarial debiasing model, using adaptive $\lambda$ policy, with $k=3$ using different optimizer setting. Metrics are shown for both pre-training and post-training stages. Columns with subscripts (1) and (2) refer to sensitive attributes- Low-Income status and First Generation flag respectively.

### v3

@@ Introduction
On training the KAN models on $D_{(1)}$ dataset, we observe the Adam-based model to offer stronger predictive power ($0.45\%$ incr. in Acc., $3.41\%$ incr. in AUROC) compared to OAdam, which enhances fairness outcomes significantly. However, the model trained using ADOPT exhibits a balanced performance across fairness and accuracy. On the other hand, training the KAN models on the larger $D_{(2)}$ dataset yields notable improvements in predictive performance and fairness. Although all setups trained on $D_{(2)}$ show significant performance improvement, the ADOPT-based model stands out by achieving balanced optimization between accuracy ($0.15\%$, and $0.82\%$ incr. in Acc. compared to Adam setup) and fairness metrics ($\{15.14\%,18.48\%\}$ incr. in $p\%$-Rule compared to Adam setup). Meanwhile, the $B_{(1)}$ and $B_{(1)}^{\prime}$ models under-perform relative to the KAN models, especially on the $D_{(2)}$ dataset. This suggests KAN-based adversarial learning to be more effective in maintaining fairness-accuracy trade-offs while capturing complex patterns, even in a feature-rich setting. SOTA baseline models, $B_{(2)}$ and $B_{(3)}$, display varying behavior. $B_{(2)}$ achieves high fairness on $D_{(1)}$ but suffers a drop on $D_{(2)}$, whereas $B_{(3)}$ starts with lower fairness on $D_{(1)}$ but improves substantially on $D_{(2)}$ with proper fine-tuning.
@@
The empirical evidence demonstrates that the KAN-based adversarial debiasing framework, regardless of the choice of optimizer, consistently outperforms the baseline models across both the datasets. This can be largely attributed to the adaptive policy usage within the training module. Its ability to dynamically adjust and bound the fairness constraint, $\lambda$, achieves to secure a balanced performance. Amongst all the configurations, the model trained on ADOPT demonstrates comparative superior balanced performance. This observation suggests ADOPT to be well suited to the spline-based architecture of KANs, owing to their adaptive convergence behavior (taniguchi2024adopt). We highlight the frameworks that demonstrate a balanced fairness-accuracy trade-off in Table 1 for clarity.
@@
Figure 2: Ablation study to compare predictive performance and fairness scores across proposed frameworks with different spline knot complexities, and baseline models under varying optimization strategies, evaluated on dataset $D_{(1)}$. (a) Illustrates metrics of the proposed KAN-based adversarial frameworks using three different optimizer techniques ($o_{1}$: Adam, $o_{2}$: OAdam, $o_{3}$: ADOPT). Similarly, (b) depicts the Baseline model performance across varying optimizers.
@@
To investigate how KAN spline knot complexity, $k$, affects the adversarial learning process, we perform an ablation study focusing on three spline orders $k=\{3,4,5\}$. We chose the $D_{(1)}$ dataset for the experiments due to its controlled small-scale setting. It enables focused assessment of the model’s sensitivity to architectural changes under constrained data conditions. In addition to the KAN-based adversarial learning models, $K_{(1)}$, we also train the baseline model, $B_{(1)}$, across the three optimizer settings, and the SOTA baseline models, $B_{(2)}$, and $B_{(3)}$ across the Adam optimizer. As illustrated in Figure 2(a), increasing the spline order $k$ leads to subtle variations in accuracy and fairness metrics. Although the $K_{(1)}$ models exhibit relatively stable accuracy across different $k$ setups, the fairness metrics, measured via the $p\%\text{-Rule}$, tend to improve with higher spline orders. This suggests that more expressive spline functions, in an adversarial setting, can better model fair representations. Although higher $k$ (e.g., $k=4,5$) may introduce convergence instability, as observed in Figure 3, the model trained with ADOPT exhibits comparatively stable training dynamics than those trained on Adam or OAdam. Next, when comparing the $K_{(1)}$ models with the baseline MLP models, $B_{(1)}$, we observe that the former consistently outperforms the latter in achieving a balance between accuracy and fairness. As observed in Figure 2(b), $B_{(1)}$ models have a higher fairness score but comparatively lower accuracy across the three optimizer settings. This further empirically substantiates the efficiency of the adaptive $\lambda$ update policy. The SOTA baseline models, $B_{(2)}$ and $B_{(3)}$, show strong performance in fairness but incur slight drops in accuracy as seen in Figure 2(b). As seen in Figure 3, the $K_{(1)}$ model training loss curve exhibits pronounced spikes. This could be attributed to the smoothness-dimension trade-off (samadi2024smooth) in KANs. While this can induce oscillatory behavior during training, it is inherent to preserving the universal approximation. A formal discussion of this property is provided in the Appendix. Despite the fluctuations in KAN-based setups, our findings help us empirically establish the efficiency of the proposed adversarial learning framework.
@@
In this study, we propose a KAN-based adversarial learning framework with adaptive $\lambda$ update policy to mitigate socioeconomic bias in a college admission setting. We discuss the various structural properties of KAN, providing proper mathematical justification behind the adversarial robustness of the networks. Through extensive experiments on two distinct real-world admissions datasets, we demonstrate the efficiency of our proposed framework by consistently outperforming SOTA baselines while achieving a balance between the inherent accuracy-fairness tradeoff. While our proposed approach effectively mitigates output-level bias, it opens up several promising enhancements. The incorporation of explicit feature-level bias detection through a reweighting mechanism could improve the overall fairness of the model while also helping to adhere to any data privacy constraints. Additionally, extending the current framework by incorporating alternative fairness definitions beyond DP could provide new insights. We believe this work offers a promising step towards aligning algorithmic performance with social responsibility, encouraging future explorations into KAN-based fairness-aware systems that could uphold both ethical and performance standards in critical decision-making.
Acknowledgments
We gratefully acknowledge the University of California, Irvine, for providing access to the admissions datasets used in this study. We also thank Prof. Iftekhar Ahmed for providing access to GPU resources used for large-scale experiments.
@@ Appendix A Additional Theoretical Analysis
As provided in Eq. 6, $f$ is composed of inner functions $\phi$ (=h(.)=h(.)), and outer functions, $\psi$, where each of them is a smooth spline. Let the inner sum be defined as , $v_{k}(x)=\sum_{i=1}^{n}\phi_{k_{i}}(x_{i})$ Then we can write $f$ as,
$f(x_{1},\dots,x_{n})=\sum_{k=1}^{2n+1}\psi_{k}(v_{k}(x))$
@@
$\left|\frac{\partial v_{k}}{\partial x_{j}}\right|=\left|\phi_{k_{j}}^{\prime}(x_{j})\right|\leq B$
@@
$\frac{\partial f}{\partial x_{j}}=\sum_{k=1}^{2n+1}\psi_{k}^{\prime}\left(v_{k}(x)\right)\cdot\phi_{k_{j}}^{\prime}(x_{j})$
@@
The input $x_{i}$ is summed at every neuron in a KAN structure, and subsequently a univariate spline $\psi(x)$ is applied. Let the sum of the inner functions be denoted as $v$. Both $\phi$, and $\psi$ are twice continuously differentiable with bounded second order derivatives. Then, we can write, $|\phi_{k_{i}}^{\prime\prime}|\leq\beta_{1}$, $|\psi_{k}^{\prime\prime}|\leq\beta_{2}$, and the derivative of inner sum $v^{\prime}(x)$ is bounded by some constant $B>0$. Using chain rule, the second derivative of the composite function $f(x)=\psi_{k}(v(x))$ can be expressed as,
$f^{\prime\prime}(x)=\psi_{k}^{\prime\prime}(v(x))\cdot[v^{\prime}(x)]^{2}+\psi_{k}^{\prime}(v(x))\cdot v^{\prime\prime}(x)$
@@
$|f^{\prime\prime}(x)|\leq\beta_{2}B^{2}+|\psi_{k}^{\prime}(v(x))|\cdot\beta_{v}$
where, $\beta_{v}$ is an upper bound on $v^{\prime\prime}(x)$. Thus each composition $\psi_{k}(\sum_{i}\phi_{k_{i}}(x_{i}))$ is $\beta$-smooth. More generally, the entire KAN, $f$, is a nested composition of such units. Hence, if each layer’s activation function is twice continuously differentiable, then the final composed function also stays $C^{2}$ smooth (deboor1978splines).
@@
In this appendix section, we discuss a supplementary special case for analyzing the fairness bounds in an adversarial debiasing setup. Although this is not a part of our main algorithmic contributions, but serves to provide additional insight into the behavior of the framework under the following conditions. We constraint the adversary model, $g$, through gradient penalty in order to satisfy and hold the following case (arjovsky2017wasserstein).
Let $(X,d_{\mathcal{X}})$ and $(Y,d_{\mathcal{Y}})$ be the metric spaces. Given, the KAN classifier function $f:X\to Y$ is $L$-Lipschitz, i.e. $d_{\mathcal{Y}}(f(x),f(y))\leq L\cdot d_{\mathcal{X}}(x,y)$ for all $x,y\in X$.
@@
For all probability distributions $\mu$ and $\nu$ on a metric space $X$, the Wasserstein-1 distance between $\mu$ and $\nu$ is defined as
$W_{1}(\mu,\nu)=\sup_{\|\varphi\|_{\mathrm{Lip}}\leq 1}\left(\mathbb{E}_{x\sim\mu}[\varphi(x)]-\mathbb{E}_{x\sim\nu}[\varphi(x)]\right),$
where the supremum is taken over all $1$-Lipschitz functions $\varphi:X\to\mathbb{R}$ (villani2008optimal).
@@
Let $\mu$, $\nu$ be two distributions on $X$, and $f:X\to Y$ be $L$-Lipschitz, then,
@@
$\leq L\,d_{\mathcal{X}}(x,y)$ (7)
Let $\varphi=\frac{1}{L}g\circ f$, or technically speaking $\varphi=\frac{1}{L}g(f(x))$. Then $\varphi$ is 1-Lipschitz on $X$ and by Definition B,
$\mathbb{E}_{\mu}[g\circ f]-\mathbb{E}_{\nu}[g\circ f]=L(\mathbb{E}_{\mu}[\varphi]-\mathbb{E}_{\nu}[\varphi])\leq LW_{1}(\mu,\nu)$
@@
$W_{1}(f_{\#}\mu,f_{\#}\nu)=\text{sup}_{\|g\|_{\text{Lip}\leq 1}}(\mathbb{E}_{\mu}[g\circ f]-\mathbb{E}_{\nu}[g\circ f])\leq LW_{1}(\mu,\nu)$
@@
Let $Z\in\{0,1\}$ be a sensitive attribute, then, $\mu:=\mathbb{P}(X|Z=0)$, $\nu:=\mathbb{P}(X|Z=1)$ are the conditional distributions of the input features $X$ given $Z=0$ and $Z=1$ respectively. For any KAN, $f$, with Lipschitz constant $L_{f}$, in order to obtain a 1-Lipschitz adversary $g$, the empirical fairness objective for the adversary model can be stated as,
$L_{adv}(g,f):=\mathbb{E}[g(f(X))|Z=0]-\mathbb{E}[g(f(X))|Z=1].$
By Proposition 3, this difference is bounded: $|L_{adv}(g,f)|\leq L_{f}\cdot W_{1}(\mu,\nu)$. Hence, maximizing $L_{adv}(g,f)$ over 1-Lipschitz $g$ yields $W_{1}(f_{\#}\mu,f_{\#}\nu)$.
@@ Appendix C Additional Experimental Results
All the experiments, excluding those involving $D_{(2)}$ datasets, have been conducted on MacBook Air with Apple M3 chip, and 24GB unified memory, running macOS. Large scale training on the $D_{(2)}$ datasets were performed on a remote Linux server with four NVIDIA RTX A6000 GPUs (49GB each), using CUDA 12.4. We have used Python-based NumPy and PyTorch packages to train the models. The hyperparameter configurations used for all experiments conducted on the proposed models are summarized in Table 2.
Table 2: Hyperparameter list for proposed KAN-based adversarial debiasing frameworks, $K_{(1)}$ (trained on $D_{(1)}$ dataset), and $K_{(2)}$ (trained on $D_{(2)}$ dataset) with an adaptive $\lambda$-update policy under three different optimizers (Adam, OAdam, and ADOPT). Width denotes the layer sizes for the classifier/adversary; $\text{LR}_{clf}$ &\& $\text{LR}_{adv}$ are learning rates; $L_{2}$ &\& $L_{1}$ penalty are regularization coefficients; $\lambda_{1}$ &\& $\lambda_{2}$ are the final fairness penalties obtained for Low Inc. status &\& First-Gen. flag sensitive features respectively; $\eta$ is the fairness learning rate.
Model Optimizer $\text{LR}_{clf}$ $\text{LR}_{adv}$ $\text{Width}_{clf}$ $\text{Width}_{adv}$ $\text{L}_{2}$ Penalty $\text{L}_{1}$ Penalty $\lambda_{(1)}$ $\lambda_{(2)}$ $\eta$
Adam 0.01 0.1 [140, 64, 32, 1] [1, 32, 2] 0.001 0.001 0.21 0.32 0.04
$\text{K}_{(1)}$ OAdam 0.01 0.1 [140, 32, 8, 1] [1, 32, 2] 0.0001 0.01 0.24 0.33 0.04
ADOPT 0.01 0.1 [140, 64, 32, 1] [1, 32, 2] 0.001 0.01 0.25 0.37 0.04
Adam 0.01 0.1 [140, 64, 8, 1] [1, 32, 2] 0.0001 0.01 0.30 0.45 0.04
$\text{K}_{(2)}$ OAdam 0.01 0.01 [140, 32, 8, 1] [1, 32, 2] 0.0001 0.01 0.31 0.43 0.04
ADOPT 0.01 0.01 [140, 32, 8, 1] [1, 32, 2] 0.0001 0.01 0.36 0.49 0.04
Justification of Optimizers
For simulations, we compare the proposed model performance across three different types of optimizers - Adam, Optimistic Adam (OAdam), and ADOPT. This comparison is motivated by the understanding that the optimizer selection significantly impacts convergence and stability in adversarial training setups (liu2020loss). Adam is included as it is a strong baseline for general NN training. Whereas, OAdam modifies Adam with an optimistic mirror-descent step, specifically designed for adversarial problems. Similarly, ADOPT tries to achieve optimal convergence rate adaptively, making it suitable for adversarial training. By conducting an ablation study we explore how different optimizers influence training dynamics and fairness outcomes.
@@
Table 3: Fairness metrics for the KAN-based adversarial debiasing model, using adaptive $\lambda$ policy, with $k=3$ using different optimizer setting. Metrics are shown for both pre-training and post-training stages. Columns with subscripts (1) and (2) refer to sensitive attributes- Low Income status and First Generation flag respectively.